OpenStudy (anonymous):

given r(x) = 6x/x+8,find r(-3)

OpenStudy (anonymous):

r*x=6x/x+8 r*x=6+8 r*x=14

OpenStudy (anonymous):

or should the right side be 6x/(x+8)?

OpenStudy (anonymous):

6(-3)/(-3+8) =-18/5